OpenStudy (anonymous):

Two angles are complementary. The sum of the measure of the first angle, and one forth of the second angle is 57 degrees. Find the measurements of the angles. Smaller and other

OpenStudy (akshay_budhkar):

you will get two equations... have you tried getting it?

OpenStudy (saifoo.khan):

complementry means, 90 - x

OpenStudy (akshay_budhkar):

saifoo let the people try it..

OpenStudy (saifoo.khan):

LOL, sure! ;)

OpenStudy (akshay_budhkar):

Cali23760 you got 2 equations?? lol saifoo

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Since they are complementary angles, their sum is 90. The second sentence gives us a+b/4=57 Thus the two equations are a+b=90 a+b/4=57

OpenStudy (akshay_budhkar):

anhhuyalex we know the answers... we must help them solve it.. not solve it for them.... we must help them learn not answer their work!

OpenStudy (anonymous):

so 90 is the smaller angle, and 57 is the other angle?

OpenStudy (anonymous):

no, the angles are a and b, you use those two equations to find what a and b are

OpenStudy (anonymous):

LOL. sorry. :|

OpenStudy (anonymous):

geez...I dont know this stuff

OpenStudy (anonymous):

we'll help you through it

OpenStudy (anonymous):

so you have your two equations, a+b=90 a+b/4=57, right?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

now we have two equations with two unknowns

OpenStudy (anonymous):

we use one equation to solve one of the unknowns in terms of the other

OpenStudy (anonymous):

i recommend subtracting b on both sides of the first equation

OpenStudy (anonymous):

so 90a?

OpenStudy (anonymous):

you should bet a=90-b

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

now this relation is also true for the second equation

OpenStudy (anonymous):

so we can plug this new value for a in the second equation

OpenStudy (anonymous):

try it and let me know what you get

OpenStudy (anonymous):

so b=57-a

OpenStudy (anonymous):

no, dont do the same thing, substitute 90-b for a in the second equation

OpenStudy (anonymous):

so a+b/4=57 becomes 90-b+b/4=57

OpenStudy (anonymous):

ok,now what

OpenStudy (anonymous):

you can solve this equation for b since it is the only variable

OpenStudy (anonymous):

ok, I am not familiar with this kind of work

OpenStudy (anonymous):

getting to this point is substitution, solving it is basic algebra

OpenStudy (anonymous):

you combine the b terms (-b+b/4=-3/4 b) and subtract 90 on both sides

OpenStudy (anonymous):

so you should end up with -3/4 b= -33

OpenStudy (anonymous):

now just multiply each side by -4/3, and you will get the answer for b

OpenStudy (anonymous):

what do you get?

OpenStudy (anonymous):

44?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

now plug that into the identity we found earlier to get a

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

we found the identity a=90-b earlier, now that we have b, we can just insert it into this equation to get a

OpenStudy (anonymous):

56

OpenStudy (anonymous):

close

OpenStudy (anonymous):

90-44=46

OpenStudy (anonymous):

i feel dumb

OpenStudy (anonymous):

its fine, i almost made the same error

OpenStudy (anonymous):

so 44 and 46

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

44+46=90, 46+44/4=46+11=57 in case you wanted to check